OpenStudy (anonymous):

Little Bill just bought a chocolate ice cream cone with his chore money. Little Bill is wondering if all the ice cream will fit inside the cone. Seeing as how he likes to experiment, Little Bill starts to push the ice cream into the cone with his fingers. Find out if all the ice cream will fit, or if it will overflow! The scoop of ice cream is a sphere with a 2.5 inch diameter. The cone has a height of 4 inches and a diameter of 2 inches. Volume of Cone=__________________ Volume of Sphere=________________ Overflow? Yes / NO

OpenStudy (ujjwal):

Volume of sphere=\(\frac{4}{3}\pi r^3\) Volume of cone= \(\frac{1}{3}\pi r^2h\) find both volumes. If volume of sphere is greater, ice-cream will overflow else not.

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

Can you just tell me what they are and then i can figure it out how you got it

OpenStudy (anonymous):

How about we call the person in the problem, Lil' Rich

OpenStudy (ujjwal):

The scoop is a sphere and the cone is a cone. Now find their volumes. r is radius and h is height in the above relations i mentioned.

OpenStudy (anonymous):

i still dont get it